# Large Language Model Can Continue Evolving From Mistakes

### v4

@@ 5.2 Multiple Iterations Analysis
We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, using the same training set construction strategy and data source as the first round. We select CIT data independent of the first round but maintain the same data volume. The experimental results presented in Appendix Table 8 show that after the second round of CEM, the accuracy of the LLM’s answers significantly improved across all dataset construction strategies. In the best case, R-Wiki shows a 4.75% improvement in accuracy compared to its previous round and a 17.00% improvement compared to INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of R-Wiki decreases by 7.02% compared to P-Wiki and by 5.90% compared to E-Wiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplementary training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM combined with the random replay method. Here, ER indicates the model’s enhancement rate on the Xiezhi task, while AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
@@
To analyze catastrophic forgetting that may arise from iterations, we use CuteGPT-13B and test it on Xiezhi, HotpotQA, and GSM8K tasks to evaluate the model’s domain knowledge, reasoning, and mathematical abilities. We employ the CEM method to enhance performance on Xiezhi and assess any performance declines on the other two tasks. Initially, we sample 2,000 instances from each of the three datasets for pre-finetuning, then test to establish baseline. We use a method of random replay to mitigate catastrophic forgetting. Specifically, in each training round, we randomly replay 500 instances from both HotpotQA and GSM8k to replace the original General Corpus.
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi still exhibits significant improvement and maintains an upward trend. Additionally, the combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70% in most cases. However, we observe that CEM-R, with its larger data capacity, achieves significant performance improvements but also shows the greatest decline on other tasks. This may be attributed to the increased volume of CPT data diluting the effect of the replay data. Combining CEM with other advanced replay-based CL methods warrants further investigation.
@@ 7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations. In the best case, accuracy enhancement reaches up to 17.00%.
Appendix A Format of CIT Data
@@
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
@@
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
@@
Question:
(question)In the 2nd century, what became the Roman Empire’s internal sea was _____
@@
(extract prompt)Please answer in the context of the following background material:
(supplement)The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history of the world.
Answer:
(answer)The answer is C. Mediterranean Sea. The complete answer is that: In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
@@
Appendix B Process of Collecting Corpora from Multiple Sources
Fig.6 and Fig.7 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
Figure 6: The process of collecting supplementary corpus from Bing.
Figure 7: The process of collecting supplementary corpus from Wikipedia.
Appendix C Experimental Results of LLMs on QA tasks
Table 3 and Table 4 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
@@ Appendix C Experimental Results of LLMs on QA tasks
Table 3: The table shows the experimental results of LLMs using the CEM method on the Xiezhi task.
Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
@@
Table 4: The table shows the experimental results of LLMs using the CEM method on the CMMLU task.
Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 55.18 54.67 53.40 49.44
W2R - 22.20 24.12 23.42 24.93 24.61 24.82 22.90
R2W - 18.57 15.02 15.65 12.68 12.87 14.35 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.86 54.53 53.90 46.58
W2R - 21.91 23.15 21.37 24.25 25.18 24.99 22.28
R2W - 18.93 17.61 19.18 13.32 13.58 14.02 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.42 54.43 53.37 53.07
W2R - 21.99 25.12 23.67 24.04 25.06 25.00 24.81
R2W - 18.31 14.16 15.81 13.55 13.46 14.56 14.67
@@ Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
Appendix E Data Division
Table 6 and Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing Supplemental Training
Xiezhi 27,265 10,000 10,000 7,265
CMMLU 11,917 4,000 5,000 2,917
Table 6: The data division and usage of Xiezhi and CMMLU.
Dataset Total Pre-Finetuning Testing Random Replay
HotpotQA 7,000 2,000 5,000 500
GSM8K 7,000 2,000 5,000 500
Table 7: The data division and usage of HotpotQA and GSM8K, where the data used for Random Replay is included in pre-finetuning.
Appendix F Experimental Results of Multiple Rounds of CEM
We employ the identical training data construction strategy as in the first round to conduct a second round of CEM on the LLM. Table 8 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after two rounds of CEM.
@@ Appendix F Experimental Results of Multiple Rounds of CEM
Appendix G Experimental Results of Catastrophic Forgetting Analysis
Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Round Strategy Task Accuracy Relative Rate
Xiezhi GSM8K HotpotQA ER AFR

### v5

@@ 5.2 Multiple Iterations Analysis
The CEM method consistently yields significant improvements across multiple rounds, underscoring its iterative potential. We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, employing the same training strategies and data sources as in the first round. Independent CIT data of equivalent volume are used. Results in Table 10 indicate that this round significantly enhances the LLM’s accuracy across all dataset construction strategies. Notably, RWiki shows a 4.75% increase in accuracy over the previous round and a 17.00% improvement over INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of RWiki decreases by 7.02% compared to PWiki and by 5.90% compared to EWiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplemental training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM. The left part represents results without replay, while the right part shows results with a random replay strategy. ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
@@
To analyze catastrophic forgetting, we test CuteGPT-13B on Xiezhi, HotpotQA, and GSM8K to evaluate domain knowledge, reasoning, and mathematical skills. We use the CEM method to improve performance on Xiezhi and monitor any declines in the other tasks. Initially, we sample 2,000 instances from each dataset to establish a baseline. To mitigate forgetting, we employ random replay by replacing the General Corpus with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70.0% in most cases. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by the random replay strategy), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 54.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods warrants further investigation.
@@ 7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
8 Limitations
This paper focuses on the model’s performance in question-answering tasks. It remains uncertain whether CEM can consistently enhance the model’s capabilities and performance in non-QA tasks.
Future research will focus on investigating additional rounds of supplemental training to further enhance the LLM and explore the integration of other strategies to mitigate catastrophic forgetting. Additionally, it will examine whether repeated supplemental learning on already mastered corpora could have adverse effects.
Furthermore, we will experiment with diverse data sources, including synthetic corpora, and optimize strategies for collecting supplementary corpora to endow the LLM with advanced continual iterative evolutionary capabilities.
Appendix A Knowledge Point Extraction
Table 1 shows the prompt used for extracting knowledge points from questions using Qwen2-7B-Instruct. In the experiment, the variable $x$ in the prompt is set to $5$.
An Example of Knowledge Point Extraction
Instruction:
(task)Please analyze the core knowledge points examined by the following question.
(question)### Question:
What is the main component of the cell membrane?
A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
(answer)### Answer
The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
(requirement)### Requirements:
- Prioritize identifying directly relevant named entities from the question and options.
- Knowledge points should be closely related to the question and options, aiding in eliminating incorrect options and selecting the correct one.
- Knowledge points should be specific, avoiding overly broad, common, or indistinct concepts, e.g., "Prime Minister", "Archaeology".
- Return the knowledge points in a list format, e.g., ["English Civil War", "Glorious Revolution"].
- The number of returned knowledge points should not exceed {x}.
Output:
(answer)["Cell Membrane Structure", "Phospholipid Bilayer"]
Table 1: An example illustrating the extraction of knowledge points from questions.
Appendix B Effectiveness Check of Supplemental Corpus
Table 5 shows the prompt used for entailment checking of the Supplemental Corpus with Qwen2-72B-Instruct. The results are presented in Table 2, where the proportion of useful information from both data sources consistently exceeds 65%.
Data Source Data Volume Effectiveness
Bing 8161 0.6718
Wiki 7694 0.6508
Table 2: The results of verifying whether the Supplemental Corpus contains information that helps the LLM in correcting errors and answering questions accurately. ‘Effectiveness’ denotes the proportion of useful information.
Appendix C Process of Collecting Corpora from Multiple Sources
The parameter configuration for the experiment is as follows: the value of $k$ is set to $4$ for top-k retrieval. Indexing is implemented using FAISS’s IndexFlatIP. The threshold for inner product similarity is set at $0.80$. If no entries meet this threshold, the entry with the highest similarity exceeding $0.70$ is selected; otherwise, no result is returned.
Fig.6 and Fig.7 show the process of obtaining supplemental data from Bing and Wikipedia respectively.
Figure 6: The process of collecting supplemental corpus from Bing.
Figure 7: The process of collecting supplemental corpus from Wikipedia.
Appendix D Format of CIT Data
@@
(question)### Question:
In the 2nd century, what became the Roman Empire’s internal sea was _____
@@
Output:
(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
@@
Instruction:
(task)Please select the correct answer for the following single choice questions based on the supplementary materials provided:
(question)### Question:
In the 2nd century, what became the Roman Empire’s internal sea was _____
@@
(supplement)### Supplementary Materials:
The Roman Empire … During the reign of Tulajin (98-117), the Roman Empire reached its greatest extent with the Mediterranean Sea becoming the empire’s inland sea. In its heyday, it controlled about 5 million square kilometers of land, making it one of the largest monarchies in the ancient history.
Output:
(answer)The answer is C. Mediterranean Sea. In the 2nd century, what became the Roman Empire’s internal sea was Mediterranean Sea.
@@
Appendix E Entailment Checking of Extractive Corpus
Table 5 shows the prompt used for entailment checking of the Extractive Corpus with Qwen2-72B-Instruct.
An Example of Entailment Checking
Instruction:
(task)Can the correct answer to the given multiple-choice question be derived solely from the knowledge in the following material? Please respond with ’Yes’ or ’No’.
(question)### Question:
What is the main component of the cell membrane?
A. Phospholipids B. Carbohydrates C. Proteins D. Nucleus
(answer)### Answer
The answer is A. Phospholipids. Phospholipids are the main component of the cell membrane.
(supplement)### Materials:
The phospholipid bilayer is a thin polar membrane made of two layers of lipid molecules. These membranes are flat sheets that form a continuous barrier around all cells. The cell membranes of almost all organisms and many viruses are made of a lipid bilayer, as are the nuclear membrane surrounding the cell nucleus, and membranes of the membrane-bound organelles in the cell…
Output:
(answer) Yes.
Table 5: An example illustrating the entailment checking of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix F Exploring Fine-Tuning Strategies
We experiment with fine-tuning strategies on Qwen-7B-Chat, employing full fine-tuning and LoRA fine-tuning for parameter updates. For dataset usage, we apply two approaches: single-stage parallel training on combined CPT and CIT data (3000 and 2000 samples, respectively), and multi-stage sequential training, where the model is first fine-tuned on CPT data, followed by further fine-tuning on CIT data.
Table 6 shows that parallel full fine-tuning yields the best results. Therefore, in the experiments, except for ChatGLM-6B, which employs Prompt-tuning Liu et al. (2022), all other models use parallel full fine-tuning.
Pre-Finetuning Parallel Sequential
Full 0.4162 0.4854 0.3154
LoRA - 0.4573 0.4271
Table 6: Results of verifying the impact of CPT and CIT datasets on model performance using different parameter update strategies and training stages.
Appendix G Data Division
Table 7 shows the the data division of datasets. For Xiezhi and CMMLU, the multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a normative format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing Supplemental Training Random Replay
Xiezhi 27,265 10,000 10,000 7,265 -
CMMLU 11,917 4,000 5,000 2,917 -
HotpotQA 7,000 2,000 5,000 - 500
GSM8K 7,000 2,000 5,000 - 500
Table 7: The data division and usage of Xiezhi, CMMLU, HotpotQA, and GSM8K. For HotpotQA and GSM8K, the data used for Random Replay is included in pre-finetuning.
Appendix H Experimental Result
Table 8 presents the comprehensive experimental results for various LLMs employing the CEM method on the Xiezhi and CMMLU tasks, with the supplemental training dataset constructed using the Plain Strategy.
Table 10 presents the comprehensive experimental results of CuteGPT-13B on the Xiezhi task following two rounds of CEM, with the second round employing the same training data construction strategy as the first.
Table 11 presents the comprehensive results of catastrophic forgetting experiments for CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Task Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
@@ Appendix H Experimental Result
CMMLU Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
@@
Table 8: Experimental results of various LLMs using the CEM method on the Xiezhi and CMMLU tasks
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0 CEM-R0.5 CEM-R1 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 53.40 55.18 54.67 49.44
W2R - 22.20 24.12 23.42 24.82 24.93 24.61 22.90
R2W - 18.57 15.02 15.65 14.35 12.68 12.87 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.90 53.86 54.53 46.58
W2R - 21.91 23.15 21.37 24.99 24.25 25.18 22.28
R2W - 18.93 17.61 19.18 14.02 13.32 13.58 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.37 53.42 54.43 53.07
W2R - 21.99 25.12 23.67 25.00 24.04 25.06 24.81
R2W - 18.31 14.16 15.81 14.56 13.55 13.46 14.67
@@
Round Strategy Xiezhi GSM8K HotpotQA ER AFR